# Measurement-induced asymmetry in bipartite networks

### v1

@@ III.2 Measurement-induced control
This section is devoted to determining whether there is an inherent advantage when one Prisoner is restricted to a given type of measurement. Initially, suppose one Prisoner is constrained to projective measurements and the other Prisoner is restricted to weak measurements.
By considering the expressions above in (12) and (13), and limiting Prisoner $B$’s measurement to be projective ($z\rightarrow\infty$ or $\tanh z\rightarrow 1$ ), one can see that this restriction does not offer any advantage, as standard game-theoretic techniques would still be required to find the optimal solution (i.e., Nash equilibrium [32]). Similar expressions would be calculated if Prisoner $A$ was restricted to projective measurements instead of Prisoner $B$.
On top of the strong measurement performed by Prisoner $B$, let us assume that Prisoner $A$ is restricted to a weak measurement ($y=0$). Thus, (12) and (13) become
@@
respectively. Now, it is seen that Prisoner $B$ can choose particular measurement directions to optimize the payoff $u_{B}$. It is also interesting to note that this depends on the correlation controlled by $x$. One solution for the maximum of $u_{B}$ would be when $\theta_{b}=\pi$ and $x=0$. This would yield a result of $u_{B}=7/4$. Therefore, it is in Prisoner $B$’s interest for Prisoner $A$ to be restricted to a weak measurement, as this allows Prisoner $B$ to maximize $u_{B}$, ensuring it is larger than $3/2$. Figure 1 shows the maximum expected payoff for both players as a function of $x$, where the maximization is taken over $\theta_{b}$ for each value of $x$ when $z\rightarrow\infty$ and $y=0$.
Figure 1: This figure shows the maximum expected payoff for both players as a function of $x$. Prisoner $B$ can choose the measurement settings ($\theta_{b}$) in order to maximize the expected payoff $u_{B}$ for each value of $x$. Therefore, this measurement-induced asymmetry gives an inherent advantage to Prisoner $B$.
@@ IV Discussion
There are a couple of key comments that are important for this work. While Bayesian games are considered to see how two parties can interact across a network, determining the Nash equilibrium has not been considered. This is because the focus of the work is revealing the measurement-induced asymmetry. It is an interesting future question to understand how this asymmetry impacts Nash equilibria and whether the measurement process allows increasingly optimal solutions.
It is clear (compare Appendix A with Appendix B) that the measurement-induced asymmetry is also strongly related to the choice of payoffs, suggesting the game-theoretic nature of this work is also important. Whether there is a general relation between the choice of payoffs and witnessing measurement-induced asymmetry remains an open question.
Moreover, while a discorded state was used to demonstrate this measurement-induced asymmetry, the state itself was symmetric, in the sense that the quantum discord is the same irrespective of which subsystem the measurement is performed on. Therefore, this further justifies the claim that the asymmetry is entirely due to the measurement process. In the future, studying how an asymmetric (one-way discorded) state affects the measurement will also be of practical relevance.
@@
The experimental realization of this work in terms of generalized measurements is a feasible task since weak measurements have been performed in superconducting qubits [48, 49]. Moreover, the bipartite network we have formulated in this work can be readily implemented in an optical experimental setup. For example, in Ref. [50], the measurement operators (4) were used.
@@ V Summary
In this paper, we have established that weak measurements naturally introduce asymmetry in a bipartite network given a discorded state when one party is restricted to projective measurements. Importantly, we have shown that this asymmetry allows a given party to control the other party’s expected payoff, as expected payoffs can be made independent of the detector settings. This could be used as a feasible technique in future quantum network design for designing systems with inherent control.
@@ Appendix A CHSH Game
Furthermore, an interesting extension of the Bell state is to consider the Werner state [52]. This is because the Werner state’s correlation evolves from classically uncorrelated, to discorded, entangled, and non-separable as a function of a parameter $\eta$. It is defined by
@@
where |$\ket{\phi}=1/\sqrt{2}\big{[}\ket{01}-\ket{10}\big{]}$.
To verify that the framework with generalized measurements produces the expected results in the projective measurement limit, it is important to consider a known example. Therefore, by using the CHSH game [53], which is a Bayesian game-theoretic version of the CHSH inequality [54], it can be established that the correct results are obtained, specifically for the Bell state. The standard CHSH game is found by considering Player $A$’s expected payoffs given in Table 2.
@@
$u_{A}^{Q}=\frac{1}{8}\Big{[}4+\Big{(}\cos(\theta_{\bm{a}}-\theta_{\bm{b}})+\cos(\theta_{\bm{a^{\prime}}}-\theta_{\bm{b}})+\cos(\theta_{\bm{a}}-\theta_{\bm{b^{\prime}}})-\cos(\theta_{\bm{a^{\prime}}}-\theta_{\bm{b^{\prime}}})\Big{)}\tanh(y)\tanh(z)\Big{]}.$ (18)
By taking $y\rightarrow\infty$, this reproduces the expected result. Furthermore, maximizing the resultant function yields $(1/8)(4+2\sqrt{2})\approx 0.85$, which is consistent with known results. It is interesting to note the maximum occurs for a projective measurement for both players. However, either player can choose to collapse the cosine terms such that $y=z=0$, which gives $u_{A}=1/2$. Since it is a zero-sum game, Player $B$’s expected payoff is the negative of Player $B$’s $(u_{Q}^{B}=-u_{A}^{Q})$.
@@
When taking the projective measurement limit, and maximizing over all variables, this gives a maximum payoff of $0.75$, which is the classical limit. This is because the discorded state is local, whereas the Bell state is non-local. Subsequently, it is clear that the expected results are reproduced for this framework. The expected payoff for the discorded state is similar in behavior to the corresponding payoff for the Bell state, in the sense that both players have the same control over the expected payoff due to their respective weak measurements. Additionally, the expected payoff for the Werner state is similar to the Bell state, except for the coupling parameter $\eta$ and a change of sign, which results in
@@ Appendix B Modified CHSH Game
Let us consider a modified combative CHSH game given in Table 3. Notice how it is still a zero-sum game, however, the payoff matrices for each type of game are no longer symmetric. It is worth emphasizing that the only modification is the payoffs.
@@
Table 3: This table shows a different payoff matrix for a modified CHSH game. This game allows inherent asymmetry to be introduced due to the measurement process. The experimental setup is the same as in Table 2. Note how the payoffs are no longer symmetric.
Using the payoffs in Table 3, and performing a similar analysis as for Table 2, the Bell state results in a similar payoff in the sense that the measurement is equivalent for both players. This implies that each player has the same amount of control based on their measurement over the expected payoff. This can be seen by
@@
Similarly, for the Werner state, the result is
@@
However, for the discorded state, there is some asymmetry introduced entirely due to the weak measurement. This confirms what is found in the main text. Explicitly, this means one player has more control over the outcome of the expected payoff than the other, as if $y\rightarrow\infty$, then the expected payoff is zero for both players. However, if $z\rightarrow 0$, the expected payoffs still have a dependence on the measurements. This can be seen in the expected payoffs given by
@@
and
$\begin{split}u_{B}^{D}&=-\frac{1}{8}\tanh(y)\Big{\{}2\big{[}\cos\theta_{\bm{a}}+\cos(\theta_{\bm{a}}-x)\big{]}\\ &+\cos\theta_{\bm{a^{\prime}}}+\cos(\theta_{\bm{a^{\prime}}}-x)-\big{[}\cos(\theta_{\bm{a^{\prime}}}-\theta_{\bm{b^{\prime}}})\\ &+\cos(\theta_{\bm{a^{\prime}}}+\theta_{\bm{b^{\prime}}}-x)\cos x\big{]}\tanh(z)\Big{\}}.\end{split}$ (24)

### v2

@@ III.2 Measurement-induced control
This section is devoted to determining whether there is an inherent advantage when one prisoner is restricted to a given type of measurement. Initially, suppose one prisoner is constrained to projective measurements and the other one is restricted to weak measurements. This type of network could be predetermined by the referee who controls the network. Given these restrictions, the prisoners would have to determine the optimal strategies in order to maximize their respective payoffs. At this stage, it is worth emphasizing that the players know the payoffs and the correlation between them so they can compute their respective expected payoffs and make optimal strategic decisions based on the known information.
By considering the expressions above in (12) and (13), and limiting prisoner $B$’s measurement to be projective ($z\rightarrow\infty$ or $\tanh z\rightarrow 1$), one can see that this restriction does not offer any advantage, in the sense that both players still have full control over their respective payoffs through measurement. Subsequently, standard game-theoretic techniques would still be required to find the optimal solution (i.e., Nash equilibrium [31]). Similar expressions would be calculated if prisoner $A$ was restricted to projective measurements instead of $B$.
Figure 1: This figure shows the maximum expected payoff for both players as a function of $x$. Prisoner $B$ can choose the measurement settings ($\theta_{b}$) in order to maximize the expected payoff $u_{B}$ for each value of $x$. Therefore, this measurement-induced asymmetry gives an inherent advantage to prisoner $B$. It is seen that the extrema occur at the classical correlation of the state, namely $x=0,n\pi$ where $n\in\mathbb{Z}$.
On top of the strong measurement performed by prisoner $B$, let us assume that prisoner $A$ is restricted to a weak measurement, which, up to a first-order series expansion around $y=0$, yields
$u_{A}(y,z;x;\theta_{a},\theta_{b},\theta_{a^{\prime}},\theta_{b^{\prime}})\equiv u_{A}=\frac{3}{2}\\ -\frac{1}{8}y\Bigg{\{}\cos\theta_{a}+\cos(\theta_{a}-x)-\cos(\frac{\theta_{b}-\theta_{b^{\prime}}}{2})\Bigg{[}\cos(\theta_{a^{\prime}}-\frac{\theta_{b}+\theta_{b^{\prime}}}{2})+\cos(\theta_{a^{\prime}}-x+\frac{\theta_{b}+\theta_{b^{\prime}}}{2})\cos x\Bigg{]}\Bigg{\}},$ (14)
and
$u_{B}(y,z;x;\theta_{a},\theta_{b},\theta_{a^{\prime}},\theta_{b^{\prime}})\equiv u_{B}=\frac{3}{2}\\ -\frac{1}{8}\Bigg{\{}\cos\theta_{b}+\cos(\theta_{b}-x)-\cos(\frac{\theta_{a}-\theta_{a^{\prime}}}{2})\Bigg{[}\cos(\theta_{b^{\prime}}-\frac{(\theta_{a}+\theta_{a^{\prime}})}{2})+\cos(\theta_{b^{\prime}}-x+\frac{\theta_{a}+\theta_{a^{\prime}}}{2})\cos x\Bigg{]}y\Bigg{\}}.$ (15)
It is clear that in the weak measurement limit, prisoner $A$ has differing levels of control over each payoff function. To explicitly highlight this, if we take the extreme case of no measurement ($y=0$), then (14) and (15) become
@@
respectively.
Now, it is seen that prisoner $B$ can choose particular measurement directions to optimize the payoff $u_{B}$. It is also interesting to note that this depends on the correlation controlled by $x$. One solution for the maximum of $u_{B}$ would be when $\theta_{b}=\pi$ and $x=0$. This would yield a result of $u_{B}=7/4$. Therefore, it is in prisoner $B$’s interest for prisoner $A$ to be restricted to a no-measurement, as this allows prisoner $B$ to maximize $u_{B}$, ensuring it is larger than $3/2$. One can envisage scenarios where the effect of making another party’s operations redundant would have significant practical benefits in the design of future interacting networks, where one party having control over the network is a necessity of the design. Subsequently, we expect measurement theory to be used as a mechanism for the future design of secure quantum networks. Figure 1 shows the maximum expected payoff for both players as a function of $x$, where the maximization is taken over $\theta_{b}$ for each value of $x$ when $z\rightarrow\infty$ and $y=0$.
@@ IV Discussion
There are a couple of key comments that are important for this work. While Bayesian games are considered to see how two parties can interact across a network, determining the Nash equilibrium was not been considered. This is because the focus of the work is revealing the measurement-induced asymmetry. It is an interesting future question to understand how this asymmetry impacts Nash equilibria and whether the measurement process allows increasingly optimal solutions.
It is clear (compare Appendix A with Appendix B) that the measurement-induced asymmetry is also strongly related to the choice of payoffs, suggesting the game-theoretic nature of this work. In fact, there is a general relation between the choice of payoffs and witnessing measurement-induced asymmetry as for an arbitrary two qubit state, the functional form of the expected payoff is the same, thus asymmetry is always present, but may only be revealed for certain choices of payoffs.
Moreover, while a discorded state was used to demonstrate the measurement-induced asymmetry in (14) and (15), the state itself was symmetric because the quantum discord is the same irrespective of which subsystem the measurement is performed on. Therefore, this further justifies the claim that the asymmetry is entirely due to the measurement process. In the future, studying how an asymmetric (one-way discorded) state affects the measurement will also be of practical relevance.
@@
The experimental realization of this work in terms of generalized measurements is a feasible task since weak measurements have been performed in, e.g., superconducting qubits [49, 50]. Moreover, the bipartite network we have formulated in this work can be readily implemented in an optical experimental setup. For example, in Ref. [51], the measurement operators (4) were used.
@@ V Summary
In this paper, we have established that generalized measurements naturally introduce asymmetry in a bipartite network given a discorded state when one party is restricted to projective measurements. Importantly, we have shown that this asymmetry allows a given party to control the other party’s expected payoff through the choice of measurements, as expected payoffs can be made independent of the detector settings. This could be used as a feasible technique in future quantum network design for designing systems with inherent control to given nodes on the network.
@@ Appendix A CHSH Game
Furthermore, an interesting extension of the Bell state is to consider the Werner state [53]. This is because the Werner state’s correlation evolves from classically uncorrelated to discorded, entangled, and non-separable as a function of a parameter $\eta$. It is defined by
@@
where |$\ket{\phi}=\dfrac{1}{\sqrt{2}}\big{[}\ket{01}-\ket{10}\big{]}$.
@@
It is important to consider a known example to verify that the framework with generalized measurements produces the expected results in the projective measurement limit. Therefore, by using the CHSH game [54], which is a Bayesian game-theoretic version of the CHSH inequality [55], it can be established that the correct results are obtained, specifically for the Bell state. The standard CHSH game is found by considering player $A$’s expected payoffs given in Table 2.
@@
$\begin{split}u_{A}^{Q}&=\frac{1}{8}\Big{[}4+\Big{(}\cos(\theta_{\bm{a}}-\theta_{\bm{b}})+\cos(\theta_{\bm{a^{\prime}}}-\theta_{\bm{b}})\\ &+\cos(\theta_{\bm{a}}-\theta_{\bm{b^{\prime}}})-\cos(\theta_{\bm{a^{\prime}}}-\theta_{\bm{b^{\prime}}})\Big{)}\\ &\times\tanh(y)\tanh(z)\Big{]}.\end{split}$ (20)
Since it is a zero-sum game, player $B$’s expected payoff is the negative of player $B$’s $(u_{Q}^{B}=-u_{A}^{Q})$. Taking $y\rightarrow\infty$ reproduces the expected result for the CHSH game [54]. Furthermore, maximizing the resultant function yields $(1/8)(4+2\sqrt{2})\approx 0.85$, which is consistent with known results. It is interesting to note the maximum occurs for a projective measurement for both players. However, either player can choose to collapse the cosine terms such that $y=z=0$ gives $u_{A}=1/2$.
@@
Taking the projective measurement limit and maximizing over all variables gives a maximum payoff of $0.75$, which is the classical limit. The discorded state is local, whereas the Bell state is non-local. Subsequently, it is clear that the expected results are reproduced for this framework. The expected payoff for the discorded state is similar in behavior to the corresponding payoff for the Bell state in the sense that both players have the same control over the expected payoff due to their respective weak measurements. Additionally, the expected payoff for the Werner state is similar to the Bell state, except for the coupling parameter $\eta$ and a change of sign, which results in
@@ Appendix B Modified CHSH Game
Let us consider a modified combative CHSH game given in Table 3. Notice how it is still a zero-sum game. However, the payoff matrices for each type of game are no longer symmetric along the diagonal elements. Therefore, there is an added complication which the players need to take into account. However, it is worth emphasizing that the only modification is the payoffs.
@@
Table 3: This table shows a different payoff matrix for a modified CHSH game. This game allows inherent asymmetry to be introduced due to the measurement process. The experimental setup is the same as in Table 2. Note how the payoffs are no longer symmetric along the diagonal elements of the individual payoff matrices.
Using the payoffs in Table 3 and performing a similar analysis as for Table 2, the Bell state results in a similar payoff because the measurement is equivalent for both players. This implies that each player has the same amount of control based on their measurement over the expected payoff. This can be seen by
@@
Similarly, for the Werner state the result is
@@
However, some asymmetry (using the definition given in Sec. III.1) is introduced entirely due to the weak measurement for the discorded state. This confirms what is found in the main text. Explicitly, this means one player has more control over the outcome of the expected payoff than the other, as if $y\rightarrow\infty$, then the expected payoff is zero for both players. However, if $z\rightarrow 0$, the expected payoffs still depend on the measurements. This can be seen in the expected payoffs given by
@@
with $u_{B}^{D}=-u_{A}^{D}$.
@@
Therefore, it is clear that the observed asymmetric effect is robust against changes in the payoffs, ensuring the measurement-induced asymmetry persists, particularly when utilizing weak measurements. This has been checked for a variety of well-known quantum states, and the observed phenomena has been confirmed in each of the cases studied.
As a final check of the robustness of the asymmetry, an arbitrary two-qubit state was taken to determine if the same functional forms found throughout this paper are reproduced for the various payoff matrices considered. It was found for each payoff matrix, the functional form was the same. Therefore, the amount of control due to the measurement is consistent, irrespective of which state is used. This further highlights the observed effects are entirely due to measurement. In particular, we have demonstrated that weak measurement can be used as a mechanism for control and advantage when parties are performing local measurements on a bipartite network.